The speed of a motor boat in still water is 60kph. It goes 100km up the river and then comes the 100km back in a total of 12 hours. what is the speed of the current of the river?
Answer by mananth(12270)

boat speed = 60 kmph
current speed = x kmph
upstream speed =60-x
downstream speed =60+x
Total distance =100km Total Time= 12.00 hours
Time upstream + time downstream =
100/(60-x)+100/(60+x)= 12
100(60+x)+100(60-x)= 12
6000+100x+6000-100x=12(15^2-x^2)
12000=12(3600-x^2)
12000 = 43200 -12 X^2
12 X ^2 = 31200
X^2=2600
x= 50.99 km/h